OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6. -112.5 -135 -90 -108 @ganeshie8 @hero @dan815 @perl @pooja195 @nincompoop @CGGURUMANJUNATH @zepedrix @usukidoll

OpenStudy (nincompoop):

use power rule

OpenStudy (anonymous):

how do i do that? can u show me step by step? @nincompoop

OpenStudy (nincompoop):

you can get the general derivative first and then apply at x=6 or you can apply it from the beginning

OpenStudy (usukidoll):

\[f(x) = -12x^2+9x \]

OpenStudy (nincompoop):

alright, guy what have you learned so far?

OpenStudy (nincompoop):

have you learned derivative by limit definition yet?

OpenStudy (anonymous):

no or maybe i did but i dont remember. can one of u show me step by step @UsukiDoll @nincompoop

OpenStudy (nincompoop):

okay do you know what a derivative is first?

OpenStudy (anonymous):

no @nincompoop

OpenStudy (nincompoop):

ahhhhh

OpenStudy (nincompoop):

do you know what a slope is?

OpenStudy (anonymous):

once i have f(x)=-12x^2+9(x) what would i do?

OpenStudy (nincompoop):

I can give you the general formula for power-rule, but it won't help you much in the long run if you have no clue what derivatives are.

OpenStudy (nincompoop):

so do you know what is slope?

OpenStudy (nincompoop):

I am asking a series of questions to see where is the proper place to begin.

OpenStudy (anonymous):

no i dont @nincompoop

OpenStudy (astrophysics):

\[\frac{ d }{ dx } x^n = nx^{n-1}\] power rule

OpenStudy (nincompoop):

man...

OpenStudy (nincompoop):

laughing out loud

OpenStudy (usukidoll):

wow nin... don't be rude -_-

OpenStudy (nincompoop):

I'd use \(a \) instead of the coefficient n

OpenStudy (nincompoop):

whatever floats floats

OpenStudy (anonymous):

@UsukiDoll can u please help me

OpenStudy (nincompoop):

alright, I think we've got a problem here I am dumbfounded why you're doing derivative when you do not know what slopes are.

OpenStudy (nincompoop):

in any case, we can go ahead and start with slope in general

OpenStudy (nincompoop):

begin by our day-to-day language, slope pertains to the steepness of something think of a slide or a hill or anything that goes down or up, it will have a slope

OpenStudy (nincompoop):

now we can try to infuse a some math into the language

OpenStudy (astrophysics):

Well I don't think you have to go that far nin, I'm sure OP knows what slope is and just needs to remember  one image will do wonders

OpenStudy (nincompoop):

I was just about to draw

OpenStudy (anonymous):

okay i already know this?

OpenStudy (nincompoop):

are you asking us or telling us?

OpenStudy (anonymous):

telling u

OpenStudy (nincompoop):

alright great! now we can move on to derivatives! laughing out loud

OpenStudy (astrophysics):

But for your question, you find the derivative then plug in the value of x they've given you

OpenStudy (usukidoll):

sorry I'm with another user right now.

OpenStudy (nincompoop):

basically derivative pertains the slope at a particular instant

OpenStudy (nincompoop):



OpenStudy (nincompoop):

a good example of the instant that I am referring to is where one of the lines intersects at one point with the circle

OpenStudy (astrophysics):

Jessica were you taught definition\[f'(x) = \lim_{h \rightarrow 0} \frac{ f(x+h)-f(x) }{ h }\] of derivative, I think that's what you have to use

OpenStudy (nincompoop):

now before we proceed, we must be able to identify and know about continuous functions

OpenStudy (nincompoop):

think of this portion as an algebra recap will you be able to tell us if the equation you're given is continuous or not?

OpenStudy (nincompoop):

if you're unsure, now is the time to learn them http://tutorial.math.lamar.edu/Classes/CalcI/CalcI.aspx

OpenStudy (nincompoop):

review on tangents with pretty awesome illustration http://tutorial.math.lamar.edu/Classes/CalcI/Tangents_Rates.aspx

OpenStudy (nincompoop):

let us know when you've digested the concept of tangents and derivatives because the next part will be testing your algebra skills with derivatives using limit-definition.